# Counting-Stars (★): A Multi-evidence, Position-aware, and Scalable Benchmark for Evaluating Long-Context Large Language Models

### v4

@@ Abstract
Despite recent efforts to develop large language models with robust long-context capabilities, the lack of long-context benchmarks means that relatively little is known about their performance. To alleviate this gap, in this paper, we propose Counting-Stars, a multi-evidence, position-aware, and scalable benchmark designed to evaluate the multi-evidence retrieval capabilities of long-context LLMs. Counting-Stars comprises two counting-based multiple pieces of evidence retrieval tasks: searching and reasoning. Using Counting-Stars, we conducted experiments to evaluate several long-context LLMs, including GPT-4 Turbo, Gemini 1.5 Pro, Claude3 Opus, GLM-4, and Moonshot-v1. Extensive experimental results demonstrate that Gemini 1.5 Pro achieves the best overall results, while GPT-4 Turbo exhibits the most stable performance across various tasks. Furthermore, our analysis of these LLMs, which have been extended to handle long-context scenarios, indicates that significant room for improvement remains as the length of the input context and the complexity of the tasks increase. The code and data are publicly accessible here [^1] https://github.com/nick7nlp/Counting-Stars.
@@ 2.1 Long-Context Multi-evidence Searching
Multi-evidence searching refers to the capability of distinguishing and collecting critical information framed within intricate and long textual data, which bottlenecks the performance of LLMs in synthesizing contextualized knowledge to execute various tasks, from answering multi-document questions to executing complex human instructions. Furthermore, maintaining a comprehensive and accurate grasp of the input text becomes increasingly challenging as the context length increases. Therefore, in the Counting-Stars test, the first task is to examine the multi-evidence searching ability of long-context LLMs, as illustrated in Table 1 (named as Long-Context Multi-evidence Searching). In multi-evidence searching, all sentences describing the little penguin counting stars are designed as "The little penguin counted {number1}". Here, {number1} indicates the number of stars the little penguin counted. Concretely, we randomly generated all the numbers of stars as {number1, number2, …} because we found that LLMs easily slack off if a sequence of numbers is continuous or regular. In this task, we hope that LLMs collect all the numbers of stars the little penguin counted and summarize them in a list.
@@ 3.1 Baselines and Experimental Settings
Generally, evaluating text-based results is usually more complex, so the evidence to be collected in the Counting-Stars is all numerical, making it more straightforward to evaluate. For a piece of test data, the prediction results are evaluated starting from the first number of stars, that is, {number1}, {number2}, …, {numberM}. In this paper, we adopt P@N as the evaluation metric, which includes two modes: one where N=M, representing the counting times (here, M denotes the counting times); the other, referencing prior studies Yuan et al. (2020); Song et al. (2023a, b), adopts the P@M as the evaluation metric, where M denotes the total number of the retrieved results.
@@ 3.2 Overall Performance
Table 7 present the testing performance of GPT-4 Turbo, Claude3 Opus, Gemini 1.5 Pro, GLM-4, Moonshot-v1 on the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Searching) and Counting-Stars-(32)-(Multi-evidence Reasoning) tests. Overall, Claude3 Opus achieves the best performance on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Searching), Gemini 1.5 Pro obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Searching) and the Chinese version of the Counting-Stars-(32)-(Multi-evidence Reasoning), and GPT-4 Turbo obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Reasoning). Although these long-context LLMs have achieved nearly perfect performance on the needle-in-a-haystack task, they still perform poorly on the Counting-Stars test, which indicates that the needle-in-a-haystack is too simple to truly show the capabilities of LLMs in processing long texts.
@@ 6 Conclusion
We introduce a multi-evidence, position-aware, and scalable benchmark for evaluating the capabilities of the multi-evidence collection of long-context LLMs via two counting-based tasks. Utilizing the Counting-Stars benchmark, we conduct intriguing analyses on the behavior of LLMs, including the length-stability dilemma and the absence of the "lost-in-the-middle" phenomenon. Our analysis provides valuable insights into how LLMs handle long contexts, which can inform and guide future research endeavors.
7 Acknowledgments
@@ 8 Limitations
While this paper offers some insights into the performance of long-context LLMs, it may not be sufficiently diverse or extensive to provide a comprehensive evaluation of the long-context capabilities of LLMs, a constraint common to most analyses and benchmarks. However, we consider that for long-context scenarios, the capability to acquire multi-evidence is the most critical capability, which is also the central aspect tested by the Counting-Stars. Finally, we analyze and summarize potential uncertainty in our experiments, experiments on more famous LLMs, and the future expansion of the Counting-Stars.
@@
(2) Experiments on more famous LLMs.
@@
Still, through our experiments, we find that even if it is a simple mathematical problem of calculating “1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1”, the probability of LLMs calculating correctly is lower. So, introducing the summation operation may be a simple and direct extension of the Counting-Stars in the future. Therefore, in the paper, we choose to let LLMs list all the numbers of stars and construct the Counting-Stars as a multi-evidence, position-aware, and scalable benchmark.
To further optimize and enhance the Counting-Stars test, we imagine other evaluation strategies similar to the Counting-Stars, such as scattering stars by different players and specifying LLMs to search for the stars counted by one of them. Based on the above idea, adding more complex interactions between players may construct a more difficult question for evaluating LLMs.

### v5

@@ Abstract
Despite recent efforts to develop large language models with robust long-context capabilities, the lack of long-context benchmarks means that relatively little is known about their performance. To alleviate this gap, in this paper, we propose Counting-Stars, a multi-evidence, position-aware, and scalable benchmark designed to evaluate the multi-evidence retrieval capabilities of long-context LLMs. Counting-Stars comprises two counting-based multiple pieces of evidence retrieval sub-tasks: searching and reasoning. Using Counting-Stars, we conduct experiments to evaluate several long-context LLMs, including GPT-4 Turbo, Gemini 1.5 Pro, Claude3 Opus, GLM-4, and Moonshot-v1. Extensive experimental results demonstrate that Gemini 1.5 Pro achieves the best overall results, while GPT-4 Turbo exhibits the most stable performance across various tasks. Furthermore, our analysis of these LLMs, which have been extended to handle long-context scenarios, indicates that significant room for improvement remains as the length of the input context and the complexity of the tasks increase. The code and data are publicly accessible here [^1] https://github.com/nick7nlp/Counting-Stars.
@@ 2.1 Long-Context Multi-evidence Searching
Multi-evidence searching refers to the capability of distinguishing and collecting critical information framed within intricate and long textual data, which bottlenecks the performance of LLMs in synthesizing contextualized knowledge to execute various tasks, from answering multi-document questions to executing complex human instructions. Furthermore, maintaining a comprehensive and accurate grasp of the input text becomes increasingly challenging as the context length increases. Therefore, in the Counting-Stars test, the first task is to examine the multi-evidence searching ability of long-context LLMs, as illustrated in Table 1 (named as Long-Context Multi-evidence Searching). In multi-evidence searching, all sentences describing the little penguin counting stars are designed as "The little penguin counted {number1}". Here, {number1} indicates the number of stars the little penguin counted. Concretely, we randomly generated all the numbers of stars as {number1, number2, …} because we found that LLMs easily slack off if a sequence of numbers is continuous or regular. In this task, we hope that LLMs collect all the numbers of stars the little penguin counted and list all digits rather than summation.
@@ 3.1 Baselines and Experimental Settings
Generally, evaluating text-based results is usually more complex, so the evidence to be collected in the Counting-Stars is all numerical, making it more straightforward to evaluate. For a piece of test data, the prediction results are evaluated starting from the first number of stars, that is, {number1}, {number2}, …, {numberM}. In this paper, we adopt P@N as the evaluation metric, which includes two modes: one where N=M, representing the counting times (here, M denotes the counting times); the other, referencing prior studies Yuan et al. (2020); Song et al. (2023a, b, 2024), adopts the P@M as the evaluation metric, where M denotes the total number of the retrieved results.
@@ 3.2 Overall Performance
Table 7 present the performance of GPT-4 Turbo, Claude3 Opus, Gemini 1.5 Pro, GLM-4, Moonshot-v1 on the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Searching) and Counting-Stars-(32)-(Multi-evidence Reasoning). Overall, Claude3 Opus achieves the best performance on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Searching), Gemini 1.5 Pro obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Searching) and the Chinese version of the Counting-Stars-(32)-(Multi-evidence Reasoning), and GPT-4 Turbo obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Reasoning). Although these LLMs have achieved nearly perfect performance on the needle-in-a-haystack task, they still perform poorly on the Counting-Stars, which indicates that the needle-in-a-haystack is too simple to truly show the capabilities of LLMs in processing long texts.
@@ 6 Conclusion
In this paper, we propose Counting-Stars, a multi-evidence, position-aware, and scalable benchmark for evaluating the multiple pieces of evidence retrieval capabilities of LLMs via two counting-based tasks. Utilizing the Counting-Stars, we conduct intriguing analyses on the behavior of LLMs, including the length-stability dilemma and the absence of the "lost-in-the-middle" phenomenon. Our analysis provides valuable insights into how LLMs handle long contexts, which can inform and guide future research endeavors.
@@ 7 Limitations
While this paper offers some insights into the performance of long-context LLMs, it may not be sufficiently diverse or extensive to provide a comprehensive evaluation of the long-context capabilities of LLMs, a constraint common to most analyses and benchmarks. However, we consider that for long-context scenarios, the capability to acquire multi-evidence is the most critical capability, which is also the central aspect tested by the Counting-Stars. Finally, we analyze and summarize potential uncertainty in our experiments, experiments on more famous LLMs.
@@
(2) Experiments on more LLMs.
@@
Still, we find that even if it is a simple mathematical problem of calculating “1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1 + 1”, the probability of LLMs calculating correctly is lower. So, introducing the summation operation may be a simple and direct extension of the Counting-Stars. However, this paper currently focuses on the multi-evidence retrieval abilities of LLMs in long contexts. Therefore, we have chosen to require LLMs to list all the numbers of stars.
To further optimize and enhance the Counting-Stars, we imagine other evaluation strategies similar to the Counting-Stars, such as scattering stars by different players and specifying LLMs to search for the stars counted by one of them. Based on the above idea, adding more complex interactions between players may construct a more difficult question for evaluating LLMs.
Acknowledgments